OpenStudy (nannie):

Correction!!! HELP!!!!!!!!!! Need to find the exact and approximate solution in three decimal places. PLease show me your work too. I am LOST! y²-16y+64=25

OpenStudy (anonymous):

factor (y-8)^2=25 take the square root of both sides y-8=5 or y-8=-5 y=13 or y=3

OpenStudy (anonymous):

@Kira, how did you do that thing? the screenshot? it's so cool

OpenStudy (nannie):

What would be the 3 decimal answer

OpenStudy (kira_yamato):

@Alex I'm using a MacBook and the application equation editor. So I type my solution in it and pressed [Command], [Shift], and [3] all at one go before I cropped it...

OpenStudy (whpalmer4):

Put in form of quadratic solution \[ax^2 + bx + c = 0\] \[y^2-16y+64-25 = 0\] \[a=1, b= -16, c= 39\] Now use solution to find values:

OpenStudy (whpalmer4):

\[y = (-b - \sqrt(b^2-4ac))/2a\]\[y = (-b + \sqrt(b^2-4ac))/2a\] Plugging in the values, we get \[y = (-(-16) + \sqrt((-16)*(-16)-4*1*39)/2 = (16 + \sqrt(256-156))/2 = 13\] \[y=(-(-16) - \sqrt((-16)*(-16)-4*1*39)/2 = (16 - \sqrt(256-156))/2 = 3\] Factoring works if you can see the factoring, but this always works, even in real life where the problems don't all have integer coefficients :-)

OpenStudy (whpalmer4):

@Kira, you can take a snapshot of just a section of your screen by doing Command-Shift-4 and sweeping out the desired rectangle

OpenStudy (anonymous):

:)

OpenStudy (nannie):

Thank you!

OpenStudy (kira_yamato):

My (or should I say our) pleasure to be of service.... :)